. The average weight of bags of maize meal packed at a particular farm is assumed to be normally distributed with a mean of 20 kg and a standard deviation of 3.0 kg. Determine the probability that a bag weighs between 19.5 kg and 22.0 kg?
1
Expert's answer
2020-04-22T14:44:18-0400

"Given \; that, μ=20, σ=3, then,\\
 P(19.5<x<22) = P(\frac{19.5-20}{3}< Z <\frac{22-20}{3})\\
=P(-0.17<Z<0.67)\\= P(0<Z<0.17)+ P(0<Z<0.67)\\
=0.0675+0.2486=0.3161\\"